In a meter bridge experiment, when a nichrome wire is in the right gap, the balancing length is 60cm. When the nichrome wire is uniformly stretched to increase its length by 20% and again connected in the right gap, the new balancing length is nearly.
Ans 51cm
1
Expert's answer
2019-04-29T09:38:29-0400

For the first situation we can simply write the equation of the meter bridge for the left and right sides:


"R_L/R_R=\frac{60}{100-60}=1.5."

When the took the nichrome wire, stretched it and connected again, its normal area "A" on the right side decreases (since the metal wire has constant volume) by factor of 1.2:


"V=Al=A_\text{stretched}\cdot1.2l,"

"A_\text{stretched}=A/1.2."

This affects the resistance in the right side:


"R_{R\text{stretched}}=\rho\frac{1.2l}{A/1.2}=1.44\rho\frac{l}{A}=1.44R_R."

For the stretched wire case we have:


"\frac{R_L}{1.44R_R}=\frac{1.5}{1.44}=\frac{x}{100-x},"

"x=51.02\text{ cm}."